OpenStudy (anonymous):

What is the molar mass of ammonia (NH3)?

OpenStudy (anonymous):

N = 14, H =1*3 bc we have 3 hydrogen ----) 14+ 3= 17 that is answer

OpenStudy (anonymous):

It's 14.007 + 3(1.008) = 17.031 g/mol if you need more accuracy.